OpenStudy (volleyballlover55):

There are (4^2)^3 ⋅ 4^0 horses on a stud farm. What is the total number of horses on the farm? 0 4^6 4^7 4^24

OpenStudy (volleyballlover55):

@jim_thompson5910

OpenStudy (volleyballlover55):

@Bobo-i-bo

OpenStudy (reemii):

what is (x^b)^c = ? (formula you have to remember)

OpenStudy (volleyballlover55):

um i dont know

OpenStudy (reemii):

it is : \((x^b)^c = x^{bc}\).

OpenStudy (volleyballlover55):

oh okk

OpenStudy (reemii):

and \(x^a \times x^b = x^{a+b}\). What is the answer then?

OpenStudy (volleyballlover55):

x^ab?

OpenStudy (reemii):

I mean, answer to your original quesiton. (the horses)

OpenStudy (volleyballlover55):

oh 4^6

OpenStudy (reemii):

right

OpenStudy (volleyballlover55):

Which number is not in scientific notation? 0.95 ⋅ 10^8 1.1 ⋅ 10^5 7.00 ⋅ 10^8 9.9 ⋅ 10^−28

OpenStudy (volleyballlover55):

@reemii a?

OpenStudy (aveline):

Correct

OpenStudy (volleyballlover55):

Which expression is rational? 6.27316543222718… square root of 2 square root of 14. square root of 49

OpenStudy (anthonyym):

Because for it to be in scientific notation the number before 10^x has to be between 1 and 10, and 0.95 is less than 1

OpenStudy (anthonyym):

Rational means that the number ends. You can see if it ends by putting it into the calculator.

OpenStudy (volleyballlover55):

D?

OpenStudy (reemii):

yes

OpenStudy (volleyballlover55):

Which of the following numbers is irrational? a fraction with numerator negative 15 and denominator 4, a fraction with numerator negative 7 and denominator 9, square root of 4, π a fraction with numerator negative 15 and denominator 4 a fraction with numerator negative 7 and denominator 9 square root of 4 π

OpenStudy (anthonyym):

I will tell you that π is an irrational number because it goes on without repeating.

OpenStudy (volleyballlover55):

@reemii is that correct? d?

OpenStudy (volleyballlover55):

@jim_thompson5910

OpenStudy (anthonyym):

Yes d is correct. -7/9 is rational because its decimal form is predictable. It's -7.9999999 repeating and the nines go on predictably.